OpenStudy (anonymous):

find the 90th term of the sequence 9,18,27...

jimthompson5910 (jim_thompson5910):

1st term: 9 2nd term: 18 3rd term: 27 So you can see that this pattern is the list of multiples of 9

jimthompson5910 (jim_thompson5910):

the first term is equal to 1*9 = 9 the second term is equal to 2*9 = 18 the third term is equal to 3*9 = 27 etc etc

OpenStudy (anonymous):

so its 810?

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

yay!:) how do i find the missing term of 5,__,27

jimthompson5910 (jim_thompson5910):

Is this an arithmetic sequence? or geometric sequence?

OpenStudy (anonymous):

arithmetic

jimthompson5910 (jim_thompson5910):

let d be the common difference

jimthompson5910 (jim_thompson5910):

so to get the next term, we add on d to get 5+d

jimthompson5910 (jim_thompson5910):

To get the next term after that, we add on another d (5+d)+d = 5 + 2d This is equal to 27 since we know this 3rd term Therefore, 5+2d = 27 d = ??

OpenStudy (anonymous):

11?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so the next term after 5 (the first term) is 5+d = 5+11 = 16

jimthompson5910 (jim_thompson5910):

Notice how adding on another 11 gives us 16+11 = 27

OpenStudy (anonymous):

how do you know if its arithmetic?

jimthompson5910 (jim_thompson5910):

you just said it was when I asked you

jimthompson5910 (jim_thompson5910):

I'm assuming they provided this information in the problem itself

jimthompson5910 (jim_thompson5910):

a sequence is arithmetic when you add the same thing to each term to get the next term

OpenStudy (anonymous):

no im on a different one now and it wants to know if its arithmetic or not

OpenStudy (the_fizicx99):

If its arithmetic you'll see that there's a common difference

jimthompson5910 (jim_thompson5910):

so for instance, the sequence 1, 3, 5, 7, 9 is arithmetic because your'e adding 2 to each term to get the next term

jimthompson5910 (jim_thompson5910):

the sequence 1, 3, 5, 7, 10 is NOT arithmetic because each term is found by adding 2 to the previous term...but jumping from 7 to 10 breaks this pattern

OpenStudy (anonymous):

okay this is a new problem... 33,37,41 whats the 19th term... i know they have a difference of 4

jimthompson5910 (jim_thompson5910):

first term: 33 common difference: 4

jimthompson5910 (jim_thompson5910):

so a1 = 33 d = 4

jimthompson5910 (jim_thompson5910):

an = a1 + d(n-1) an = 33 + 4(n-1) an = 33 + 4n-4 an = 4n + 29

jimthompson5910 (jim_thompson5910):

So the nth term is an = 4n + 29

jimthompson5910 (jim_thompson5910):

Notice for say term 2, n = 2 an = 4n + 29 a2 = 4(2) + 29 a2 = 8 + 29 a2 = 37 which matches up with the list

OpenStudy (anonymous):

i got 105:)

OpenStudy (the_fizicx99):

There's a faster way instead of doing all of that. an = a1 + d(n – 1) you want to find the 19th term so you put a_19 = 33 + 4(19-1) a_19 = 33 + 4(18) a_19 = 33 + 72 a_19 = 105

jimthompson5910 (jim_thompson5910):

105 is correct

OpenStudy (anonymous):

what about this one?

OpenStudy (anonymous):

nevermind i got 5

jimthompson5910 (jim_thompson5910):

5 is correct, nice work